### v1

# A discussion on the symmetry of relativistic Vlasov gas and its accretion in Kerr-Newman black hole

Abstract
We investigate the kinetic properties of collisionless Vlasov gas in Kerr-Newman spacetime, focusing on how spacetime symmetries constrain the distribution functions. By establishing the general relationship between distribution functions and constants of motion (mass $m$, energy $E$, angular momentum $L_{z}$, Carter constant $L$) along with $L$’s conjugate configuration variable $Q^{3}$, we derive the analytical framework for particle dynamics in stationary axisymmetric spacetimes. Within the Locally Non-Rotating Frame , we compute key observables—particle number density, energy density, principal pressures, and accretion rates—resolving coordinate divergences near the event horizon through geometric rescaling. For a plasma obeying the Jüttner distribution, asymptotic expressions at spatial infinity are explicitly calculated. Numerical analyses reveal: (1) mass and energy accretion rates exhibit nearly identical parametric dependencies on black hole parameters; (2) extremal Kerr-Newman black holes show monotonic enhancement in mass, energy, and angular momentum accretion rates with increasing $a$, contrasting with Kerr black hole behavior. These results provide new insights into accretion mechanisms in extreme spacetime geometries.
I Introduction
Research on accretion can be traced back to the early work of Lyttleton, Hoyle and Bondi, who first provided a relativistic solution for perfect fluids Hoyle(1939) ; Bondi(1944) ; Bondi(1952) , and these were later extended to general relativity by Michel Michel(1972) . When particle collisions become infrequent, the perfect fluid approximation breaks down. In such collisionless regimes, kinetic theory governed by the Vlasov equation provides a physically viable alternative.
The theory of gas dynamics was first introduced to special relativity by Jüttner Juttner(1911) , and was later extended by Synge, Tauber, Israel, and others, and gradually incorporated into general relativity Synge(1934) ; Tauber(1961) ; Israel(1963) ; Ehlers(1971) ; Ehlers(1973) ; Cercignani(2002) . The study of kinetic theory of gases helps to understand the properties of accretion disks and the jet launching mechanism, such as reconstructing images of supermassive black holes at the center of M87 galaxy Narayan(1995) . In addition, observations indicate that the gas surrounding M87 and Sgr A* is nearly collisionless and magnetized Akiyama(2019) ; Akiyama(2022) . For studies on collisionless gas models in galaxies and galaxy clusters, see Refs. Rein(2014) ; Hadvzic(2021) .
Recent years have witnessed extensive investigation of relativistic Vlasov gases. Sarbach and his collaborators established a foundational paradigm, first analyzing relativistic kinetic gases and their distribution functions, and then constructing a rigorous geometric framework on both tangent and cotangent bundles Sarbach(201303) ; Sarbach(201309) ; Sarbach(201311) ; Sarbach(2022) . Rioseco implemented this framework numerically in Schwarzschild spacetime, revealing stark contrasts between Vlasov gas accretion and perfect fluid accretion. These results partially explain the observed accretion rates substantially below Bondi predictions Rioseco-Sarbach(20171) ; Rioseco-Sarbach(20172) . Gamboa Gamboa-Sarbach(2021) and Liao Liao(2022) incorporated angular momentum distributions, addressing phenomena like low luminosity. Mach et al. further modeled accretion onto moving Schwarzschild black holes using Vlasov gases for dark matter simulations, achieving robust results via Monte Carlo methods Mach(20211) ; Mach(20212) ; Mach(2022) ; Mach(2023) ; Mach(2024) .
For non-Schwarzschild black holes, Cieślik Cieslik-Mach(2020) and Li Li(2025) demonstrated that black hole charge suppresses mass accretion rates, and considered both charged black holes and charged gases. In stationary axisymmetric spacetimes, they developed Kerr accretion models for gas confined to the equatorial plane Cieslik-Mach(2022) ; Mach-Odrzywolek(2023) and unbound configurations Li(2023) . Discussions on the phase space structure of gas in stationary axisymmetric spacetimes and related applications can be found in Refs. Rioseco-Sarbach(2023) ; Rein(2014) ; Rein(2023) ; Amesl(2016) ; Cai(2023) .
This paper is organized as follows. Sec. II discusses how the Vlasov equation and spacetime symmetries constrain the distribution function in Kerr-Newman (KN) spacetime. Sec. III analyzes charged particle trajectories. Sec. IV derives physical quantities: particle number density, pressures, accretion rates, and asymptotic behavior. Sec. V presents numerically computed profiles of these physical quantities. Finally, Sec. VI offers a concluding summary of the work.
We adopt these conventions: a 4-dimensional Lorentzian manifold is denoted by $(M,g)$, with local coordinates $x^{\mu}$, Greek letters $\mu,\nu,...\in\{0,1,2,3\}$. The metric signature is (,+,+,+)(-,+,+,+)( - , + , + , + ), with geometric units $G=c=1$. The tangent bundle is $TM$, with $T_{x}M$ being the tangent space at $x\in M$. The cotangent bundle is $T^{*}M$, with $T^{*}_{x}M$ the cotangent space at $x$. Lie derivatives along vector field $v$ are denoted $\mathscr{L}_{v}$.
II Vlasov gas and its distribution function in Kerr-Newman spacetime
II.1 One-Particle Distribution Function and Its Invariance
In spacetime $(M,g)$, the dynamical state of a particle with mass $m$ is specified by a position $x\in M$ and physical momentum $p_{a}\in T^{*}_{x}M$. The future mass shell is
$\Gamma^{+}_{m}:=\left\{(x,p_{a})\in T^{*}M:x\in M,p_{a}\in F^{+}{x}(m)\right\},$ (1)
where $F^{+}_{x}(m)$ is the future mass hyperboloid:
[…]
The one-particle distribution function is defined as a mapping
$f:\Gamma^{+}\to\mathbb{R},$ (3)
where $\Gamma^{+}\equiv\displaystyle\cup_{m>0}\Gamma^{+}_{m}$. The particle current density $J_{\mu}$ and energy-momentum tensor $T_{\mu\nu}$ are
$J_{\mu}|_{x}$ $:=\int_{F^{+}_{x}}p_{\mu}f\mathrm{dvol}_{x}(p),$ $T_{\mu\nu}|_{x}$ $:=\int_{F^{+}_{x}}p_{\mu}p_{\nu}f\mathrm{dvol}_{x}(p),$ (4)
where $\mathrm{dvol}_{x}(p)$ represents the volume element on $F^{+}_{x}\equiv\displaystyle\cup_{m>0}F^{+}_{x}(m)$.
For free particles (influenced solely by the spacetime background) with mass $m$ and charge $q$ in Kerr-Newman (KN) spacetime with electromagnetic field $F_{ab}=(\mathrm{d}A)_{ab}$, the equations of motion for a trajectory $C(\tau)$ are
$T^{c}\nabla_{c}T^{a}=\frac{q}{m}F^{a}_{\ \,b}T^{b},$ (5)
where $T\equiv\frac{\partial\ }{\partial\tau}$ denotes the tangent vector with affine parameter $\tau$, $F^{a}_{\ \,b}\equiv g^{ac}F_{cb}$ and $A_{a}$ is the gauge potential. The trajectory $C(\tau)\equiv\left\{x^{\mu}(\tau)\right\}$ induces a curve on $T^{*}M$,
$\hat{C}(\tau):=\left\{(x^{\mu}(\tau),p_{\nu}(\tau)):p_{\mu}(\tau)\equiv mg_{\mu\nu}|_{x(\tau)}T^{\nu}(\tau)\right\},$ (6)
where $T^{\mu}\equiv\frac{\mathrm{d}x^{\mu}}{\mathrm{d}\tau}$. Since $f$ serves as the one-particle distribution function, it satisfies
$\mathscr{L}_{\hat{L}}f=\frac{\mathrm{d}f\circ\hat{C}(\tau)}{\mathrm{d}\tau}=0,$ (7)
where $\hat{L}$ denotes the tangent vector field to the curve $\hat{C}(\tau)$.
Next we consider Killing vector fields on $M$. Let $\phi:M\to M$ be the isometry generated by a Killing vector field $K$. The original trajectory $C(\tau)$ naturally induces a transformed trajectory $\tilde{C}(\tau)\equiv\phi\circ C(\tau)$, with tangent vector $\tilde{T}^{c}|_{\phi(x)}=\phi_{*}[T^{c}|_{x}]$, which preserves the dynamical equations due to the invariance properties,
$\tilde{T}^{c}\tilde{\nabla}_{c}\tilde{T}^{a}|_{\phi(x)}=\phi_{*}[(T^{c}\nabla_{c}T^{a})|_{x}]=\left.\left(\frac{q}{m}F^{a}_{\ \,b}\tilde{T}^{b}\right)\right|_{\phi(x)}.$ (8)
This derivation employs the invariance of $g$ and $\nabla$ under the pushforward/pullback operations induced by $\phi$ (see Lee(2018) ), and the electromagnetic field’s invariance under isometries $\mathscr{L}_{K}A=0$ (equivalently $\phi_{*}[A|_{x}]=A|_{\phi(x)}$).
For a Killing vector field $K^{a}$ generating an isometry group $\varPhi\equiv\left\{\phi_{k}\right\}$, we obtain a family of free particles $C(k,\tau)\equiv\phi_{k}\circ C(\tau)$. The spacetime symmetry allows the identification of these trajectories. The invariance of $f$ on the surface $\hat{C}(k,\tau)$ yields
$\mathscr{L}_{\hat{K}}f=0,$ (9)
where $\hat{K}\equiv\frac{\partial\ }{\partial k}$ denotes the tangent vector to the parameterized surface $\hat{C}(k,\tau)$.
Beyond these symmetry considerations, we impose no additional constraints on the distribution function $f:\Gamma^{+}\to\mathbb{R}$.
II.2 Distribution Function in The Phase Space
The KN metric in Boyer-Lindquist (BL) coordinates $\{t,r,\theta,\varphi\}$ is
$\mathrm{d}s^{2}=$ $-\frac{\varDelta-a^{2}\sin^{2}\theta}{\rho^{2}}\mathrm{d}t^{2}-\frac{2a\sin^{2}\theta(r^{2}+a^{2}-\varDelta)}{\rho^{2}}\mathrm{d}t\mathrm{d}\varphi$
$+\frac{\varSigma^{2}}{\rho^{2}}\sin^{2}\theta\mathrm{d}\varphi^{2}+\frac{\rho^{2}}{\varDelta}\mathrm{d}r^{2}+\rho^{2}\mathrm{d}\theta^{2},$ (10)
where $\rho^{2}=r^{2}+a^{2}\cos^{2}\theta,\varDelta=r^{2}-2\mathcal{M}r+a^{2}+Q^{2},\varSigma^{2}=(r^{2}+a^{2})^{2}-\varDelta a^{2}\sin^{2}\theta$, and $a=\frac{J}{M}$. The electromagnetic vector potential is given by
$A_{a}\equiv-\frac{Qr}{\rho^{2}}\left[(\mathrm{d}t)_{a}-a\sin^{2}\theta(\mathrm{d}\varphi)_{a}\right].$ (11)
Killing vectors $\xi^{a}=\left(\frac{\partial\ }{\partial t}\right)^{a},\psi^{a}=\left(\frac{\partial\ }{\partial\varphi}\right)^{a}$ and Killing tensorsCarter(1973) ; Wald(1984) $g_{ab},K_{ab}\equiv 2\rho^{2}l_{(a}k_{b)}+r^{2}g_{ab}$ (where $l^{a}\equiv\frac{1}{\ \Delta\ }\left[(r^{2}+a^{2})\xi^{a}+a\psi^{a}+\Delta\left(\frac{\partial\ }{\partial r}\right)^{a}\right],k^{a}\equiv\frac{1}{2\rho^{2}}\left[(r^{2}+a^{2})\xi^{a}+a\psi^{a}-\Delta\left(\frac{\partial\ }{\partial r}\right)^{a}\right]$ ) provide constants of motion
$E$ $:=-\pi_{a}\xi^{a},$ $L_{z}$ $:=\pi_{a}\psi^{a},$ $H$ $:=\frac{1}{2}g_{ab}p^{a}p^{b},$ $L^{2}$ $:=K_{ab}p^{a}p^{b},$ (12)
where $p^{a}\equiv mT^{a}$ is the physical momentum and $\pi^{a}\equiv p^{a}+qA^{a}$ is the canonical momentum.
Using the these constants of motion as the canonical momentum coordinates,
$P_{0}$ $\equiv-\sqrt{-2H},$ $P_{1}$ $\equiv E,$ $P_{2}$ $\equiv L_{z},$ $P_{3}$ $\equiv L=\sqrt{L^{2}},$ (13)
the conjugate configuration variables are defined via
$Q^{\mu}:=\frac{\partial S}{\partial P_{\mu}},$ (14)
where $S=\frac{1}{2}m^{2}\tau+\int\pi_{\mu}(x^{\alpha},P_{\beta})\mathrm{d}x^{\mu}$ denotes the action functional for free particles Straumann(2013) .
In Appendix A, we prove that the Hamiltonian vector fields satisfy
$Z^{A}_{H}$ $=m\hat{L}^{A}=g^{\mu\nu}p_{\nu}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{A}-\frac{1}{2}p_{\alpha}p_{\beta}\frac{\partial g^{\alpha\beta}}{\partial x^{\mu}}\left(\frac{\partial\ }{\partial p_{\mu}}\right)^{A}-qp_{\rho}F^{\rho}_{\ \,\sigma}\left(\frac{\partial\ }{\partial p_{\sigma}}\right)^{A},$ (15a)
$Z^{A}_{\pi(K)}$ $=\hat{K}^{A},$ (15b)
where $\pi(K)\equiv\left[\pi_{\mu}(\mathrm{d}x^{\mu})_{A}\right]\left[K^{\mu}\left(\frac{\partial\ }{\partial x}^{\mu}\right)^{A}\right]$ and $K^{a}\equiv K^{\mu}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{a}$ is a Killing vector on $M$. This yields the explicit form of the Vlasov equation Eq. (7):
$0=m\mathscr{L}_{\hat{L}}f$ $=g^{\mu\nu}p_{\nu}\frac{\partial f}{\partial x^{\mu}}-\frac{1}{2}p_{\alpha}p_{\beta}\frac{\partial g^{\alpha\beta}}{\partial x^{\mu}}\frac{\partial f}{\partial p_{\mu}}-qp_{\rho}F^{\rho}_{\ \,\sigma}\frac{\partial f}{\partial p_{\sigma}}.$ (16)
In the canonical coordinates $\left\{Q^{\alpha},P_{\beta}\right\}$, we have
$Z_{H}$ $=-P_{0}\frac{\partial\ }{\partial Q^{0}},$ $-Z^{A}_{\pi(\xi)}$ $=Z_{E}=\frac{\partial\ }{\partial Q^{1}},$ $Z^{A}_{\pi(\psi)}$ $=Z_{L_{z}}=\frac{\partial\ }{\partial Q^{2}}.$ (17)
[…]
$\frac{\partial f}{\partial Q^{0}}$ $=0,$ $\frac{\partial f}{\partial Q^{1}}$ $=0,$ $\frac{\partial f}{\partial Q^{2}}$ $=0.$ (18)
Thus, in the canonical coordinates $\left\{Q^{\alpha},P_{\beta}\right\}$, the distribution has the general form
$f=f(Q^{3},P_{0},P_{1},P_{2},P_{3}),$ (19)
We currently have no justification to eliminate additional variables from $f$ beyond the constraints derived above.
III Particle Trajectory Analysis
The equations of motion Eq. (12) for charged particles in KN spacetime are
$(\rho^{2}\dot{r})^{2}$ $=\left[(r^{2}+a^{2})E-aL_{z}-\kappa r\right]^{2}-\varDelta(L^{2}+r^{2}),$ (20a)
$(\rho^{2}\dot{\theta})^{2}$ $=L^{2}-a^{2}\cos^{2}\theta-\left(\frac{L_{z}}{\sin\theta}-aE\sin\theta\right)^{2},$ (20b)
$\rho^{2}\dot{t}$ $=\left(r^{2}+a^{2}\right)\cdot\frac{(r^{2}+a^{2})E-aL_{z}-\kappa r}{\varDelta}+a\left(L_{z}-aE\sin^{2}\theta\right),$ (20c)
$\rho^{2}\dot{\varphi}$ $=a\cdot\frac{(r^{2}+a^{2})E-aL_{z}-\kappa r}{\varDelta}+\sin^{-2}\theta\left(L_{z}-aE\sin^{2}\theta\right),$ (20d)
where $\kappa\equiv qQ$. To facilitate subsequent discussions, we now define several dimensionless parameters:
$a$ $\leftarrow\frac{a}{M},$ $Q$ $\leftarrow\frac{Q}{M},$ $x^{\mu}$ $\leftarrow\frac{x^{\mu}}{M},$ $\tau$ $\leftarrow\frac{\tau}{M},$ (21)
$E$ $\leftarrow\frac{E}{m},$ $L_{z}$ $\leftarrow\frac{L_{z}}{mM},$ $L^{2}$ $\leftarrow\frac{L^{2}}{(mM)^{2}},$ $q$ $\leftarrow\frac{q}{m}.$ (22)
The angular motion Eq. (20b) yields the following parametrized system
$p_{\theta}$ $=\bar{L}\cos\sigma,$ $L_{z}$ $=aE\sin^{2}\theta+\bar{L}\sin\sigma\sin\theta.$ (23)
where $\bar{L}\equiv\sqrt{L^{2}-a^{2}\cos^{2}\theta}$. The radial motion equation (20a) can be reformulated as
$(\rho^{2}\dot{r})^{2}=R$ $\equiv\left((r^{2}+a^{2})E-aL_{z}\right)^{2}-\varDelta(L^{2}+r^{2})$
$=\left(\rho^{2}E-\mathcal{A}\bar{L}-\kappa r\right)^{2}-\varDelta(\bar{L}^{2}+\rho^{2}),$ (24)
where $\mathcal{A}\equiv a\sin\theta\sin\sigma$. Physically, real motion requires $R\geq 0$. We only consider particles whose trajectory can extend to infinity, which requires
$\displaystyle\mathop{\mathrm{lim}}_{r\to\infty}(\dot{r})^{2}=\displaystyle\mathop{\mathrm{lim}}_{r\to\infty}\frac{R(r)}{\rho^{4}}=E^{2}-1\geq 0,$ (25)
so, $E\geq 1$ is necessary.
At a given spatial position $(r,\theta)$, the angular motion Eq. (20b) requires $\bar{L}\geq 0$, while the radial motion Eq. (24) demands $R(r)\geq 0$. These constraints define the parameter ranges $D_{\text{abs}}$ for particles absorbed by the black hole and $D_{\text{scat}}$ for scattered particles:
$D_{\text{abs}}$ $\equiv D_{\sigma}\times D_{E\text{abs}}\times D_{\bar{L}\text{abs}}=[0,2\pi]\times[1,+\infty)\times(0,\bar{L}_{c}),$ (26a)
$D_{\text{scat}}$ $\equiv D_{\sigma}\times D_{E\text{scat}}\times D_{\bar{L}\text{scat}}=[0,2\pi]\times[E_{c},+\infty)\times(\bar{L}_{c},\bar{L}_{\text{max}}),$ (26b)
where the parameter $\bar{L}$ (rather than $L$) is adopted due to its convenience when studying specific spatial positions. The upper bound $\bar{L}_{\mathrm{max}}=\bar{L}_{\mathrm{max}}(r,\theta;\sigma,E)$ is determined by the $R(r,\theta;\sigma,E,\bar{L})=0$,
$\bar{L}_{\mathrm{max}}=\frac{1}{\varDelta-\mathcal{A}^{2}}\cdot\left[-\mathcal{A}(\rho^{2}E-\kappa r)+\sqrt{\varDelta\left[(\rho^{2}E-\kappa r)^{2}-\rho^{2}(\varDelta-\mathcal{A}^{2})\right]}\right].$ (27)
The critical parameters $E_{c}-r$ and $\bar{L}_{c}-E$ can be determined by solving the periastron equation ($\dot{r}=0$): $R=0$ and $R^{\prime}\equiv\frac{\partial R}{\partial r}=0$. Simplifying under the condition $02rE-\kappa>$ 0 yields:
$2E$ $=\frac{\kappa}{r}+\sqrt{\frac{\varDelta}{\bar{L}^{2}+\rho^{2}}}+\left(1-\frac{1}{r}\right)\sqrt{\frac{\bar{L}^{2}+\rho^{2}}{\varDelta}},$ $\rho^{2}E$ $=\kappa r+\sqrt{\varDelta(\bar{L}^{2}+\rho^{2})}-\mathcal{A}\bar{L},$ (28)
Decoupling $E$ and $\bar{L}$ (involving roots of quartic equations) gives formal expressions $E=f_{E}(a,Q,r,\theta,\sigma,q)$, and $\bar{L}=f_{\bar{L}}(a,Q,r,\theta,\sigma,q)$. Numerical analysis reveals two critical radii for $r>r_{H}=1+\sqrt{1-(a^{2}+Q^{2})}$: a singularity $r_{\text{ph}}=r_{\text{ph}}(a,Q,\theta,\sigma)$ where $f_{E}\to+\infty$ and a root $r_{\text{mb}}=r_{\text{mb}}(a,Q,\theta,\sigma,q)$ where $f_{E}=1$. The monotonicity of $f_{E}-r$ over $(r_{\text{ph}},r_{\text{mb}})$ permits the formal derivation of $\bar{L}=g_{\bar{L}}(a,Q,\theta,E,\sigma,q)$
$E_{c}$ =$=\begin{cases}+\infty&\text{for }r\leq r_{\text{ph}};\\ f_{E}(a,Q,r,\theta,\sigma,q),&\text{for }r_{\text{ph}}<r<r_{\text{mb}};\\ 1,&\text{for }r\geq r_{\text{mb}}.\end{cases},$ $\bar{L}_{c}$ =$=\begin{cases}g_{\bar{L}}(a,Q,\theta,E,\sigma,q)&\text{for }E>1;\\ \bar{L}_{\text{mb}}&\text{for }E=1.\end{cases}$ (29)
where $\bar{L}_{\text{mb}}\equiv f_{\bar{L}}|_{r=r_{\text{mb}}}$. Numerical examples are shown in Fig. 1, while analytical derivations of $E_{c}$ and $\bar{L}_{c}$ are detailed in Appendix B.
(a) $E_{c}-r$
(b) $L_{c}-E_{c}$
Figure 1: For a KN black hole ($M=1$, $a=0.1$, $Q=0.1$), critical parameters (a) $E_{c}$ at the periastron radii $r$ and (b) $L_{c}-E$ on the equatorial plane ($\theta=\pi/2$) for various $\kappa$, with $\sigma=\pi/2$.
Fig. 1a shows $E_{c}$ at periastron radii. For fixed $\kappa$, $E_{c}(r)$ has a vertical asymptote at $r=r_{\text{ph}}$ and $E_{c}(r_{\text{mb}})=1$. When $r\in(r_{\text{ph}},r_{\text{mb}})$, $E_{c}$ increases as $r$ decreases. No scattering occurs for $r<r_{\text{ph}}$, while any energy $E_{c}\geq 1$ permits scattering when $r>r_{\text{mb}}$. Fig. 1b shows that $L_{c}$ grows approximately linearly with $E$ for fixed $\kappa$, with curves shifting downward as $\kappa$ increases.
IV Physical Observables of the Vlasov Gas System
IV.1 Observable Quantities
The particle flux density $J_{\mu}$ and the energy-momentum tensor $T_{\mu\nu}$ for Vlasov gases are defined by Eq. (4). We first explicitly formulate the phase space integration. The volume element
$\mathrm{dvol}_{x}(p)=\sqrt{|\mathrm{det}[g^{\mu\nu}]|}\mathrm{d}p_{t}\wedge\mathrm{d}p_{r}\wedge\mathrm{d}p_{\theta}\wedge\mathrm{d}p_{\varphi}=\frac{L}{\sqrt{R}}\mathrm{d}E\wedge\mathrm{d}m\wedge\mathrm{d}L\wedge\mathrm{d}\sigma.$ (30)
Thus,
$J_{\mu}$ $=\int_{F^{+}_{x}}p_{\mu}\bar{f}_{4}\frac{\bar{L}}{\sqrt{R}}\mathrm{d}\sigma\mathrm{d}E\mathrm{d}\bar{L},$ $T_{\mu\nu}$ $=\int_{F^{+}_{x}}p_{\mu}p_{\nu}\bar{f}_{5}\frac{\bar{L}}{\sqrt{R}}\mathrm{d}\sigma\mathrm{d}E\mathrm{d}\bar{L},$ (31)
utilizing $L\mathrm{d}L=\bar{L}\mathrm{d}\bar{L}$ and $\bar{f}_{n}\equiv\int^{+\infty}_{0}m^{n}f\mathrm{d}m$, where $p_{\mu}\leftarrow\frac{p_{\mu}}{m},E\leftarrow\frac{E}{m}$ and $\bar{L}\leftarrow\frac{\bar{L}}{m}$, hereafter denote quantities per unit mass. The momentum components are
$p_{t}$ $=-E+\frac{\kappa r}{\rho^{2}},$ $p_{r}$ =±$=\pm\frac{\sqrt{R}}{\varDelta},$ $p_{\theta}$ $=\bar{L}\cos\sigma,$ $p_{\varphi}$ $=a\sin^{2}\theta\left(E-\frac{\kappa r}{\rho^{2}}\right)+\bar{L}\sin\theta\sin\sigma.$ (32)
The integrals over absorption and scattering domains are
$J^{\mathrm{abs}}_{\mu}$ $=\mathop{\int}_{D_{\mathrm{abs}}}p_{\mu}\bar{f}_{4}\cdot\frac{\bar{L}}{\sqrt{R}}\mathrm{d}\sigma\mathrm{d}E\mathrm{d}\bar{L},$ $J^{\mathrm{scat}}_{\mu}$ $=2\mathop{\int}_{D_{\mathrm{scat}}}p_{\mu}\bar{f}_{4}\cdot\frac{\bar{L}}{\sqrt{R}}\mathrm{d}\sigma\mathrm{d}E\mathrm{d}\bar{L},$ (33a)
$T^{\mathrm{abs}}_{\mu\nu}$ $=\mathop{\int}_{D_{\mathrm{abs}}}p_{\mu}p_{\nu}\bar{f}_{5}\cdot\frac{\bar{L}}{\sqrt{R}}\mathrm{d}\sigma\mathrm{d}E\mathrm{d}\bar{L},$ $T^{\mathrm{scat}}_{\mu\nu}$ $=2\mathop{\int}_{D_{\mathrm{scat}}}p_{\mu}p_{\nu}\bar{f}_{5}\cdot\frac{\bar{L}}{\sqrt{R}}\mathrm{d}\sigma\mathrm{d}E\mathrm{d}\bar{L},$ (33b)
where the factor 2 in scattering terms accounts for symmetric pre-/post-scattering trajectories.
The divergences satisfy (see Theorem 4 in Sarbach(2022) )
$\nabla^{a}J_{a}$ $=0,$ $\nabla^{a}T_{ab}=qF_{bc}J^{c}.$ (34)
The current density $\mathcal{J}^{a}=qJ^{a}$ interacts with electromagnetic fields via $\nabla^{a}F_{ab}=-4\pi\mathcal{J}_{b}$. The electromagnetic energy-momentum tensor $\mathscr{T}_{ab}=\frac{1}{4\pi}\left(F_{ac}F^{\ c}_{b}-\frac{1}{4}g_{ab}F_{de}F^{de}\right)$ satisfies $\nabla^{a}\mathscr{T}_{ab}=-F_{bc}\mathcal{J}^{b}$. The total energy-momentum tensor
$\mathcal{T}_{ab}$ $=T_{ab}+\mathscr{T}_{ab},$ $\nabla^{a}\mathcal{T}_{ab}$ $=0,$ (35)
remains divergence-free.
For multi-species Vlasov gases (classified by charge), each $f^{\alpha}$ satisfies the Vlasov-Maxwell Eq. (16), with total current $J^{a}=\sum_{\alpha}J^{a}_{\alpha}$.
The particle number density in the comoving frame is
$n_{J}:=\sqrt{-g^{ab}J_{a}J_{b}}.$ (36)
All subsequent analysis adopts the Locally Non-Rotating Frame (LNRF) with tetrad Bardeen(1972) ,
$(e_{0})^{a}$ $=e^{-\nu}\left[(\partial_{t})^{a}+\omega(\partial_{\varphi})^{a}\right],$ $(e_{3})^{a}$ $=e^{-\varPsi}(\partial_{\varphi})^{a},$ $(e_{1})^{a}$ $=e^{-\mu_{1}}(\partial_{r})^{a},$ $(e_{2})^{a}$ $=e^{-\mu_{2}}(\partial_{\theta})^{a},$ (37a)
$(e^{3})_{a}$ $=e^{\varPsi}\left[(\mathrm{d}\varphi)_{a}-\omega(\mathrm{d}t)_{a}\right],$ $(e^{0})_{a}$ $=e^{\nu}(\mathrm{d}t)_{a},$ $(e^{1})_{a}$ $=e^{\mu_{1}}(\mathrm{d}r)_{a},$ $(e^{2})_{a}$ $=e^{\mu_{2}}(\mathrm{d}\theta)_{a},$ (37b)
where $\partial_{x}\equiv\frac{\partial\ }{\partial x},e^{2\nu}=\frac{\rho^{2}\varDelta}{\varSigma^{2}},e^{2\varPsi}=\frac{\varSigma^{2}\sin^{2}\theta}{\rho^{2}},e^{2\mu_{1}}=\frac{\rho^{2}}{\varDelta},e^{2\mu_{2}}=\rho^{2},\omega=\frac{2Mar-Q^{2}}{\varSigma^{2}}$. The particle number density becomes
$n=-J_{a}(e_{0})^{a}=-e^{-\nu}(J_{t}+\omega J_{\varphi}),$ (38)
and the energy density is given by (Note that the electromagnetic fields are not considered here)
$\varepsilon:=T_{ab}(e_{0})^{a}(e_{0})^{b}=e^{-2\nu}(T_{tt}+2\omega T_{t\varphi}+\omega^{2}T_{\varphi\varphi}).$ (39)
The principal pressures correspond to the eigenvalues of the stress tensor (the spatial projection of the energy-momentum tensor)
$\hat{T}^{i}_{\ \ j}=T^{a}_{\ \ b}(e^{i})_{a}(e_{j})^{b}=\begin{bmatrix}e^{-2\mu_{1}}T_{rr}&0&e^{-\mu_{1}-\varPsi}T_{r\varphi}\\ 0&e^{-2\mu_{2}}T_{\theta\theta}&0\\ e^{-\mu_{1}-\varPsi}T_{\varphi r}&0&e^{-2\varPsi}T_{\varphi\varphi}\end{bmatrix}.$ (40)
The principal pressures are
[…]
The divergence-free energy-momentum tensor $\nabla^{a}\mathcal{T}_{ab}=0$ permits the construction of conserved currents $\varXi_{a}:=-\mathcal{T}_{ab}\xi^{b},\varPsi_{a}:=\mathcal{T}_{ab}\psi^{b}$. The mass, energy, and angular momentum accretion rates of the black hole are respectively defined as
$\dot{\mathscr{M}}$ $:=\int_{S}J_{a}\bar{n}^{a},$$\dot{\mathscr{E}}$ $:=\int_{S}\varXi_{a}\bar{n}^{a},$$\dot{\mathscr{L}}$ $:=\int_{S}\varPsi_{a}\bar{n}^{a},$ (42)
where $S$ denotes a 2-surface enclosing the black hole and $\bar{n}^{a}$ is the outward unit normal.
IV.2 Asymptotic Analysis
We consider a simple plasma Vlasov gas system consisting of two particle species with charges $+q$ and $-q$ but otherwise identical properties, maintaining overall electrical neutrality. The distribution function and the particle current density can be compactly expressed as
$f$ $=\varrho^{+}f^{+}+\varrho^{-}f^{-},$ $J$ $\equiv\int pf\mathrm{dvol}_{x}(p)=J^{+}+J^{-},$ $\cdots$,⋯ , (43)
with $\varrho^{\pm}=\frac{1}{2}$. We assume that monospecies charged gases (i.e., gases composed of identical particles carrying charges of identical sign and magnitude) obey the simplest form of the Jüttner distribution (details in Rioseco-Sarbach(20171) ):
$f^{+}=f^{-}=\alpha_{0}\delta(P_{0}-m)e^{-zE},$ (44)
where $\alpha_{0}$ is the normalization coefficient, $m$ is the particle mass, $z\equiv\frac{mc^{2}}{k_{B}T}$ and $E\leftarrow\frac{E}{mc^{2}}$. The integral expressions for $J_{\mu}$ and $T_{\mu\nu}$ become:
$[J^{\pm}]^{\mathrm{abs}}_{\mu}$ $=\frac{1}{2}[\alpha_{0}m^{4}]\int^{2\pi}_{0}\mathrm{d}\sigma\int^{+\infty}_{1}\mathrm{d}E\int^{\bar{L}_{c}}_{0}\mathrm{d}\bar{L}\cdot\frac{\bar{L}e^{-zE}}{\sqrt{R}}\cdot p_{\mu},$ (45a)
$[J^{\pm}]^{\mathrm{scat}}_{\mu}$ $=[\alpha_{0}m^{4}]\int^{2\pi}_{0}\mathrm{d}\sigma\int^{+\infty}_{E_{c}}\mathrm{d}E\int^{\bar{L}_{\mathrm{max}}}_{\bar{L}_{c}}\mathrm{d}\bar{L}\cdot\frac{\bar{L}e^{-zE}}{\sqrt{R}}\cdot p_{\mu},$ (45b)
$[T^{\pm}]^{\mathrm{abs}}_{\mu\nu}$ $=\frac{1}{2}[\alpha_{0}m^{5}]\int^{2\pi}_{0}\mathrm{d}\sigma\int^{+\infty}_{1}\mathrm{d}E\int^{\bar{L}_{c}}_{0}\mathrm{d}\bar{L}\cdot\frac{\bar{L}e^{-zE}}{\sqrt{R}}\cdot p_{\mu}p_{\nu},$ (45c)
$[T^{\pm}]^{\mathrm{scat}}_{\mu\nu}$ $=[\alpha_{0}m^{5}]\int^{2\pi}_{0}\mathrm{d}\sigma\int^{+\infty}_{E_{c}}\mathrm{d}\bar{L}\int^{\bar{L}_{\mathrm{max}}}_{\bar{L}_{c}}\mathrm{d}\bar{L}\cdot\frac{\bar{L}e^{-zE}}{\sqrt{R}}\cdot p_{\mu}p_{\nu},$ (45d)
As $r\to+\infty$, asymptotic results become $\theta$-independent (to $\mathscr{O}(1)$ precision),
$[J^{\pm}]^{\mathrm{scat}}_{t}/[\alpha_{0}m^{4}]$ $=-\frac{2\pi K_{2}(z)}{z},\qquad\qquad[J^{\pm}]^{\mathrm{scat}}_{r}/[\alpha_{0}m^{4}]=\frac{2\pi(1+z)}{z^{3}e^{z}},$ (46a)
$[T^{\pm}]^{\mu\ \mathrm{sct}}_{\ \nu}/[\alpha_{0}m^{5}]$ $=2\pi\cdot\operatorname{diag}\left[-\left(\frac{K_{1}(z)}{z}+\frac{3K_{2}(z)}{z^{2}}\right),\frac{K_{2}(z)}{z^{2}},\frac{K_{2}(z)}{z^{2}},\frac{K_{2}(z)}{z^{2}}\right],$ (46b)
where $K_{n}(z)$ denotes the modified Bessel functions of the second kind. All the other components asymptotically vanish. The asymptotic results of some physical quantities are
$n_{J\infty}$ $=4\pi\sqrt{\left(\frac{K_{2}(z)}{z}\right)^{2}-\left(\frac{1+z}{z^{3}e^{z}}\right)^{2}},$ $n_{\infty}$ $=4\pi\frac{K_{2}(z)}{z},$ (47a)
$\varepsilon_{\infty}$ $=4\pi\left(\frac{K_{1}(z)}{z}+\frac{3K_{2}(z)}{z^{2}}\right),$ $\mathcal{P}_{i}$ $=4\pi\frac{K_{2}(z)}{z^{2}},\quad\text{for}\ i=1,2,3.$ (47b)
Near the event horizon ($r\to r_{H}$), the normalization factors $e^{-\nu}$ and $e^{-\mu_{1}}$) of the LNRF tetrad $(e_{0})^{a}$ and $(e_{1})^{a}$ induce divergences. To address this, we systematically remove these divergent factors, resulting in rescaled quantities that retain physical significance exclusively near the horizon. These modified quantities explicitly isolate finite horizon-scale physics from the divergent coordinate artifacts inherent to the LNRF tetrad formulation.
The physical quantities involved in divergence and their corrections are as follows:
$n$ $\sim e^{-\nu},$ $\varepsilon$ $\sim e^{-2\nu},$ $\mathcal{P}_{1}$ $\sim e^{2\mu_{1}},$ (48a)
$\tilde{n}$ $\equiv e^{\nu}\cdot n,$ $\tilde{\varepsilon}$ $\equiv e^{2\nu}\cdot\varepsilon,$ $\tilde{\mathcal{P}}_{1}$ $\equiv e^{-2\mu_{1}}\cdot\mathcal{P}_{1}.$ (48b)
When the surface $S$ in Eq. (42) approaches the event horizon $S\to\mathcal{B}$, and $\bar{n}^{a}\to(e_{1})^{a}$, the accretion rates simplify to:
$\dot{\mathscr{M}}$ $=S_{\mathcal{B}}\cdot\int^{\frac{\pi}{2}}_{0}\mathrm{d}\theta\sin\theta\cdot\left(e^{\mu_{1}}J^{r}\right),$ (49a)
$\dot{\mathscr{E}}$ $=S_{\mathcal{B}}\cdot\int^{\frac{\pi}{2}}_{0}\mathrm{d}\theta\sin\theta\cdot\left(-e^{\mu_{1}}\mathcal{T}^{r}_{\ \ t}\right),$ (49b)
$\dot{\mathscr{L}}$ $=S_{\mathcal{B}}\cdot\int^{\frac{\pi}{2}}_{0}\mathrm{d}\theta\sin\theta\cdot\left(e^{\mu_{1}}\mathcal{T}^{r}_{\ \ \varphi}\right),$ (49c)
where the prefactor $S_{\mathcal{B}}\equiv 4\pi(r^{2}_{H}+a^{2})$ represents the area of the event horizon $\mathcal{B}$. Considering the associated divergent factor $\frac{1}{\sqrt{\varDelta}}$ ( $e^{\mu_{1}}=\sqrt{\rho^{2}/\varDelta}$), we introduce auxiliary parameters to analyze the accretion rate $\dot{\mathcal{A}}(a,Q)$ more effectively:
$\dot{\mathcal{A}}_{\varDelta}$ $\equiv\displaystyle\mathop{\mathrm{lim}}_{r\to r_{H}}\sqrt{\varDelta}\dot{\mathcal{A}},$ $\dot{\mathcal{A}}_{\delta}$ $\equiv\frac{\dot{\mathcal{A}}_{\varDelta}}{S_{\mathcal{B}}}.$ (50)
$\dot{\mathcal{A}}_{\delta}$ can be physically interpreted as the mean accretion rate per unit area.
V Numerical Results
This section presents numerical implementations of the physical quantities discussed in Sec. III. For gas properties: $M=1$, $a=Q=0.1$, $\kappa=0,\pm 0.3$, $z=1$. For accretion rates: $a,Q\in[0,1]$. The Vlasov gas and black hole form a coupled system; $Q$ may vary but $\kappa$ is fixed.
(a) $n_{J}-\ln(r/r_{H})$
(b) $n-\ln(r/r_{H})$
Figure 2: (a) The particle number density $n_{J}/n_{J\infty}$ of the gas’s comoving frame and (b) the particle number density density $n/n_{\infty}$ of the LNRF on the equatorial plane. Dashed lines in (b) show rescaled $\tilde{n}$.
Fig. 2a shows that on the equatorial plane, $n_{J}$ decreases with $r$. At fixed $r$, the plasma and the neutral gas densities nearly coincide; the negatively charged gas is enhanced, and the positively charged gas is suppressed. Fig. 2b shows $n$ decreasing with $r$. The tetrad factor $e^{-\nu}$ causes $n$ to diverge near $r_{H}$; removing this yields finite $\tilde{n}$ (dashed).
Plasma observables closely match those of neutral gas. Thus, the subsequent plots label only the plasma curves; the neutral gas profiles are shown with black dotted lines.
(a) $\varepsilon-\log(r/r_{H})$
(b) $\mathcal{P}-\log(r/r_{H})$
Figure 3: (a) The energy density $\varepsilon$ and the rescaled energy $\tilde{\varepsilon}\equiv e^{-2\nu}\varepsilon$. (b) The principal pressures $\mathcal{P}_{i}$ and the rescaled radial pressure $\tilde{\mathcal{P}}_{1}\equiv e^{-2\mu_{1}}\mathcal{P}_{1}$.
Fig. 3a shows $\varepsilon$ decreasing monotonically with $r$. Fig. 3b shows that the pressure components decrease with $r$. The pressure anisotropy ($\mathcal{P}_{1}\neq\mathcal{P}_{2}\simeq\mathcal{P}_{3}$) and the suppression of $\tilde{\mathcal{P}}_{1}$ near $r_{H}$ confirm that a Vlasov gas differs from isotropic fluids.
The following employs numerical methods to investigate the accretion rates of Kerr black holes, RN black holes, and extreme KN ($a^{2}+Q^{2}\to 1$) black holes, focusing on their dependence on the rotation parameter $a$ and electric charge $Q$. Specific accretion rates are denoted as
$\dot{\mathcal{A}}_{S}$ $\equiv\dot{\mathcal{A}}(0,0),$ Kerr $\equiv\dot{\mathcal{A}}(a,0),$ $\text{KN}_{a}$ $\equiv\dot{\mathcal{A}}\left(a,\sqrt{1-a^{2}}\right),$
RN $\equiv\dot{\mathcal{A}}(0,1-Q),$ $\text{KN}_{Q}$ $\equiv\dot{\mathcal{A}}\left(\sqrt{1-Q^{2}},1-Q\right),$ (51)
where parameters $a$ and $Q$ vary within $[0,1]$. In the subsequent plots, the abscissa represents $x=a$ for both the Kerr and KNa curves, whereas for the RN and KNQ curves, the abscissa corresponds to $x=1-Q$.
(a) ˙$\dot{\mathcal{M}}_{\varDelta}-x$
(b) ˙$\dot{\mathcal{M}}_{\delta}-x$
Figure 4: The mass accretion rate parameters of characteristic black holes relative to the Schwarzschild case are presented.
(a) ˙$\dot{\mathcal{E}}_{\varDelta}-x$
(b) ˙$\dot{\mathcal{E}}_{\delta}-x$
Figure 5: The energy accretion rate parameters of characteristic black holes relative to the Schwarzschild case are presented.
Figs. 4 and 5 present the mass and energy accretion rates of characteristic black holes normalized by their Schwarzschild counterparts. Numerical results demonstrate identical curves between mass accretion parameters (˙$\left(\frac{\dot{\mathcal{M}}_{\varDelta}}{\dot{\mathcal{M}}_{\varDelta S}},\frac{\dot{\mathcal{M}}_{\delta}}{\dot{\mathcal{M}}_{\delta S}}\right)$ and energy accretion parameters (˙$\left(\frac{\dot{\mathcal{E}}_{\varDelta}}{\dot{\mathcal{E}}_{\varDelta S}},\frac{\dot{\mathcal{E}}_{\delta}}{\dot{\mathcal{E}}_{\delta S}}\right)$. The integral expressions for the mass and energy accretion rate parameters become nearly identical upon substituting $J^{r}$ and $T^{r}_{\ \ t}$, owing to the exponential decay of the distribution function in particle energy $E\in[1,\infty)$. This equivalence originates from specific features of our accretion model.
Figs. 4a and 4b demonstrate distinct rotational dependencies of Kerr black hole accretion rates: the mass accretion rate ˙$\dot{\mathcal{M}}_{\varDelta}$ decreases with growing rotation parameter $a$, while the average accretion rate per unit area ˙$\dot{\mathcal{M}}_{\delta}$ increases. These results align with the theoretical frameworks in Cieslik-Mach(2022) and Li(2023) respectively. The mass accretion ˙$\dot{\mathcal{M}}_{\varDelta}$ of RN black holes decreases with increasing charge parameter $Q$. However, the accretion rate parameter ˙$\dot{\mathcal{M}}_{\delta}$ exhibits distinct behaviors: it follows a non-monotonic trend with $Q$—initially decreasing before rising over the full parameter range. Extreme Kerr-Newman black holes exhibit contrasting behavior — both accretion rate parameters increase monotonically with $a$ (equivalently decreasing $Q$).
(a) ˙$-\dot{\mathcal{L}}_{\varDelta}-x$
(b) ˙$-\dot{\mathcal{L}}_{\delta}-x$
Figure 6: The angular momentum accretion rate parameters of characteristic black holes relative to the particle number density $n_{J\infty}$ are presented.
Fig. 6 shows angular momentum accretion vanishes for non-rotating black holes ($a=0$). For Kerr and extremal KN, the angular momentum accretion rate parameters ˙$-\dot{\mathcal{L}}_{\Delta}$ and ˙$-\dot{\mathcal{L}}_{\delta}$ increase monotonically with $a$. Unlike mass/energy rates, Kerr’s angular momentum curve lies inside extremal KN’s loop-shaped curves.
VI Conclusion
In this paper, we have studied the properties and specific applications of the Vlasov gas distribution function in Kerr spacetime. The main conclusions are as follows.
We have analyzed how spacetime symmetries constrain one-particle distribution functions. In the KN spacetime, $f$ for free particles depends only on the constants of motion $m,E,L_{z},L$, and the conjugate variable $Q^{3}$.
In the LNRF, we have computed the particle number density $n$, the energy density $\varepsilon$, the principal pressures $\mathcal{P}_{i}$, and the accretion rates, resolving horizon divergences via rescaling. For a plasma governed by the Jẗtner distribution, we have explicitly calculated the asymptotic expressions at spatial infinity. Numerically, all these quantities decrease monotonically with $r$, matching Schwarzschild asymptotics. At finite $r$, the plasma and the neutral gas observables nearly coincide; the positively (negatively) charged gas’s components are suppressed (enhanced).
These characteristic black holes reveal nearly identical curves between mass accretion rate parameters (˙$(\dot{\mathcal{M}}_{\varDelta},\dot{\mathcal{M}}_{\delta})$ and energy accretion rate parameters (˙$(\dot{\mathcal{E}}_{\varDelta},\dot{\mathcal{E}}_{\delta})$. An increase in both the rotation parameter $a$ and charge parameter $Q$ suppresses the mass accretion rate ˙$\dot{\mathcal{M}}_{\varDelta}$ (equivalently, the energy accretion rate ˙$\dot{\mathcal{E}}_{\varDelta}$). In contrast, the angular momentum accretion rate ˙$\dot{\mathcal{L}}_{\varDelta}$ exhibits different dependencies: it increases with rising $a$ but decreases with increasing $Q$.
Acknowledgments
This work was supported by the Young Natural Science Foundation of Wuhan Donghu College. I sincerely thank my colleague Yu-qin Yao for computational assistance in generating the numerical data for $J_{\mu}$ and $T_{\mu\nu}$ through execution of predefined computational workflows.
Appendix A The Hamiltonian Vector Fields
The tangent vector to the curve $\hat{C}(\tau)$ is given by
$L^{A}=\left(\frac{\partial\ }{\partial\tau}\right)^{A}=\frac{\mathrm{d}x^{\mu}}{\mathrm{d}\tau}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{A}+\frac{\mathrm{d}p_{\nu}}{\mathrm{d}\tau}\left(\frac{\partial\ }{\partial p_{\nu}}\right)^{A},$ (52)
where
$m\frac{\mathrm{d}p_{\nu}(\tau)}{\mathrm{d}\tau}=\frac{1}{2}p^{\rho}p^{\sigma}\partial_{\nu}g_{\rho\sigma}+qF_{\nu\gamma}p^{\gamma}=-\frac{1}{2}p_{\alpha}p_{\beta}\frac{\partial g^{\alpha\beta}}{\partial x^{\nu}}-qp_{\gamma}F^{\gamma}_{\ \,\nu}.$
This derivation employs the equations of motion Eq. (5). The tangent vector $L^{A}$ thus becomes
$mL^{A}=g^{\mu\sigma}p_{\sigma}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{A}-\left[\frac{1}{2}p_{\alpha}p_{\beta}\frac{\partial g^{\alpha\beta}}{\partial x^{\nu}}+qp_{\gamma}F^{\gamma}_{\ \,\nu}\right]\left(\frac{\partial\ }{\partial p_{\nu}}\right)^{A}.$ (53)
For the Killing vector field $K^{a}$ on $M$ with its induced one-parameter isometry group $\varPhi\equiv\left\{\phi_{k}\right\}$, the identity $\mathscr{L}_{K}K=[K,K]=0$ yields $\phi_{*}[K|_{x}]=K|_{\phi(x)}$. When $x$ lies on a particle trajectory $C(\tau)$, $\phi_{k}(x)$ resides on trajectory $C(k,\tau)$. This implies
$(\phi_{k})^{*}[(\tilde{K}^{c}\tilde{\pi}_{c})|_{\phi_{k}(x)}]$ $=K^{c}|_{x}(\phi_{k})^{*}[(\tilde{p}_{c}+qA_{c})|_{\phi_{k}(x)}]=K^{c}|_{x}[p_{c}|_{x}+(\phi_{-k})_{*}[A_{c}|_{\phi(x)}]]$
$=K^{c}|_{x}(p_{c}+qA_{c})|_{x}=(K^{c}\pi_{c})|_{x}.$ (54)
Thus one has
$0=\mathscr{L}_{K}(\tilde{K}^{c}\tilde{\pi}_{c})=\left.\frac{\mathrm{d}\ }{\mathrm{d}k}\right|_{k=0}(K^{c}\tilde{\pi}_{c})=K^{\nu}\left[p_{\alpha}\frac{\partial K^{\alpha}}{\partial x^{\nu}}+\left.\frac{\mathrm{d}\tilde{p}_{\nu}}{\mathrm{d}k}\right|_{k=0}\right].$ (55)
Consequently, the tangent vector $\hat{K}\equiv\frac{\partial\ }{\partial k}$ to $\hat{C}(k,\tau)$ becomes
$\hat{K}^{A}\equiv\left(\frac{\partial\ }{\partial k}\right)^{A}$ $=\frac{\mathrm{d}x^{\mu}}{\mathrm{d}k}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{A}+\frac{\mathrm{d}\tilde{p}_{\nu}}{\mathrm{d}k}\left(\frac{\partial\ }{\partial p_{\nu}}\right)^{A}$
$=K^{\mu}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{A}-p_{\alpha}\frac{\partial K^{\alpha}}{\partial x^{\nu}}\left(\frac{\partial\ }{\partial p_{\nu}}\right)^{A},$ (56)
On the cotangent bundle $T^{*}M\equiv\left\{(x^{\mu},p_{\nu})\right\}$, the vector fields are expressed as
$Z^{A}\equiv X^{\alpha}\left(\frac{\partial\ }{\partial x^{\alpha}}\right)^{A}+Y_{\beta}\left(\frac{\partial\ }{\partial p_{\beta}}\right)^{A}.$ (57)
The symplectic form $\varOmega_{AB}$, defined as the exterior derivative of the Poincaré 1-form $\Theta_{A}\equiv p_{\mu}(\mathrm{d}x^{\mu})_{A}$, satisfies $\varOmega_{AB}=(\mathrm{d}p_{\mu})_{A}\wedge(\mathrm{d}x^{\mu})_{B}$. For charged particles with canonical momentum $\pi_{a}\equiv\pi_{\mu}(\mathrm{d}x^{\mu})_{a}$, the symplectic form in conjugate coordinates $\left\{x^{\mu},\pi_{\nu}\right\}$ becomes (seeAcuna(2022))
$\varOmega_{AB}\equiv(\mathrm{d}\pi_{\mu})_{A}\wedge(\mathrm{d}x^{\mu})_{B}=(\mathrm{d}p_{\mu})_{A}\wedge(\mathrm{d}x^{\mu})_{B}+q\hat{F}_{AB},$ (58)
where $\hat{F}_{AB}=\frac{1}{2}F_{\mu\nu}(\mathrm{d}x^{\mu})_{A}\wedge(\mathrm{d}x^{\nu})_{B}$. The Hamiltonian vector field $Z_{F}^{A}$ for a function $\mathcal{F}$ satisfies
$\mathrm{d}\mathcal{F}=-\mathrm{i}_{Z_{H}}\varOmega,$ (59)
where the interior product is defined as
$\mathrm{i}_{Z}\varOmega:=Z^{A}\varOmega_{AB}=-X^{\alpha}(\mathrm{d}p_{\alpha})_{B}+(Y_{\beta}+qX^{\gamma}F_{\gamma\beta})(\mathrm{d}x^{\beta})_{A}.$ (60)
The Hamiltonian vector fields for these functions $H=\frac{1}{2}g^{\alpha\beta}p_{\alpha}p_{\beta}$ and $\pi(K)\equiv K^{\gamma}\pi_{\gamma}$ are respectively
$Z^{A}_{H}$ $=g^{\mu\sigma}p_{\sigma}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{A}-\left[\frac{1}{2}p_{\alpha}p_{\beta}\frac{\partial g^{\alpha\beta}}{\partial x^{\nu}}+qp_{\gamma}F^{\gamma}_{\ \,\nu}\right]\left(\frac{\partial\ }{\partial p_{\nu}}\right)^{A}=mL^{A},$ (61a)
$Z^{A}_{\pi(K)}$ $=K^{\mu}\left(\frac{\partial\ }{\partial x^{\mu}}\right)^{A}-p_{\alpha}\frac{\partial K^{\alpha}}{\partial x^{\nu}}\left(\frac{\partial\ }{\partial p_{\nu}}\right)^{A}=\hat{K}^{A}.$ (61b)
In the canonical coordinates $\left\{Q^{\alpha},P_{\beta}\right\}$, the symplectic form becomes
$\varOmega_{AB}=\left(\frac{\partial\pi_{\alpha}}{\partial P_{\mu}}\right)\mathrm{d}P_{\mu}\wedge\mathrm{d}x^{\alpha}=\frac{\partial^{2}S\ }{\partial P_{\mu}\partial x^{\alpha}}\mathrm{d}P_{\mu}\wedge\mathrm{d}x^{\alpha}=\frac{\partial Q^{\mu}}{\partial x^{\alpha}}\mathrm{d}P_{\mu}\wedge\mathrm{d}x^{\alpha}=\mathrm{d}P_{\mu}\wedge\mathrm{d}Q^{\mu},$ (62)
which yields the Hamiltonian vector fields for $H,E=-\pi(\xi)$ and $L_{z}=\pi(\psi)$,
$Z^{A}_{H}$ $=-P_{0}\left(\frac{\partial\ }{\partial Q^{0}}\right)^{A},$ $Z^{A}_{E}$ $=\left(\frac{\partial\ }{\partial Q^{1}}\right)^{A}=-Z^{A}_{\pi(\xi)},$ $Z^{A}_{L_{z}}$ $=\left(\frac{\partial\ }{\partial Q^{2}}\right)^{A}=Z^{A}_{\pi(\psi)}.$ (63)
Appendix B Periastron Analysis
For $\kappa=0$, the decoupled solutions of Eq. (28) become:
$E_{0}$ $=\frac{1}{2}\left(\sqrt{\varDelta B}+\left(1-\frac{1}{r}\right)\frac{1}{\sqrt{\varDelta B}}\right),$ $\bar{L}_{0}$ $=\sqrt{\frac{1}{B}-\rho^{2}},$ (64)
where $B=A-2C^{2}\rho^{2}-2C\sqrt{1-A\rho^{2}+C^{2}\rho^{4}}$ with $A\equiv\frac{2}{\rho^{2}}-\frac{r-1}{r\varDelta}$ and $C\equiv-\frac{a\sin\theta\sin\sigma}{\rho^{2}\sqrt{\varDelta}}$. These parameters satisfy $A=B+2C\sqrt{1-\rho^{2}B}$. As $B\to 0$ (implying $E_{0}\to\infty,\bar{L}_{0}\to\infty$), we find $A=2C$, leading to the equation for $r_{\text{ph}}$:
$-2a\sin\theta\sin\sigma=\frac{1}{\sqrt{\varDelta}}\left(2\varDelta-\rho^{2}\left(1-\frac{1}{r}\right)\right).$ (65)
For $\kappa\neq 0$, decoupling involves solving quartics, but the $r_{\text{ph}}$ equation from $E\to\infty$, $\bar{\mathcal{L}}\to\infty$ matches the $\kappa=0$ case.
On the equator ($\theta=\pi/2$), the photon radius simplifies to:
$-2a\sin\sigma=\frac{r^{2}-3r+2(a^{2}+Q^{2})}{\sqrt{r^{2}-2r+(a^{2}+Q^{2})}}.$ (66)
For $a=0$,
$r^{2}-3r+2Q^{2}$ $=0,$ $r_{\mathrm{ph}}$ $=\frac{3}{2}\left(1+\sqrt{1-\frac{8Q^{2}}{9}}\right),$ (67)
reducing to $r_{\mathrm{ph}}=3$ for the Schwarzschild black hole. For the extreme KN black holes $(a^{2}+Q^{2}\to 1,r_{H}\to 1$):
$-2a\sin\sigma$ $=\frac{r^{2}-3r+2}{\sqrt{r^{2}-2r+1}}=r-2,$ $r_{\mathrm{ph}}$ $=2-2a\sin\sigma.$ (68)
Equatorial bound orbits $p_{\theta}=0\Rightarrow\sigma=\pm\frac{\pi}{2}$ yield prograde ($L_{z}>0$) and retrograde ($L_{z}<0$) photon radii:
$r_{\mathrm{ph}-p}$ $=2-2a,$ $r_{\mathrm{ph}-r}$ $=2+2a.$ (69)
For small $\kappa$, perturbative solutions exist. Reformulating Eq. (28) in $\mathcal{E}\equiv\frac{1}{E},\mathcal{L}\equiv\frac{1}{\bar{L}}$:
$E$ $\simeq\frac{1}{\mathcal{E}_{0}+\kappa\mathcal{E}_{1}},$ $\bar{L}$ $\simeq\frac{1}{\mathcal{L}_{0}+\kappa\mathcal{L}_{1}},$ (70)
where $\mathcal{E}_{0}=1/E_{0}$, and $\mathcal{L}_{0}=1/\bar{L}_{0}$ for $\kappa=0$. The first-order corrections follow from $\mathscr{O}(\kappa)$ terms.